You are pushing an 800 kg car down the road, and it is going at 6 m/s west.
How much momentum does it have?
1
Expert's answer
2019-12-13T10:07:53-0500

The magnitude of momentum:


"p=mv=(800)(6)=4800\frac{kg\cdot m}{s}"

The direction is westwards.